# A Point-Based Algorithm for Distributional Reinforcement Learning in Partially Observable Domains

### v1

@@ 5.1 Environments
This environment, adapted from the MiniGruid suite (Chevalier-Boisvert et al., 2023, MiniGrid-DoorKey-5x5-v0), places an agent in a 5x5 grid containing a locked door and a key. The agent’s objective is to reach a goal state behind a locked door requiring a key. We assume both perfect sensor and transition models, a discount factor $\gamma=0.9$, a reward of $1.0$ upon reaching the goal state, and a known start state. Because the start state is known and the dynamics models are perfect, we set $\mathcal{B}$ to include complete confidence in being in one of the ten states along the optimal path (i.e., $|\mathcal{B}|=10$).
@@
5.2 Setup
We disable belief-point expansion in both environments and fix a finite set $\mathcal{B}$ of belief states as described in Section 5.1. PBVI is implemented as in (Pineau et al., 2003), while DPBVI follows the backup procedure from Section 4. Return distributions in DPBVI use 51 categorical atoms, with each $\psi$-vector initialized to place all probability mass at zero. We fix the convergence criterion to $\epsilon=1\mathrm{e}{-3}$.
@@ 5.2 Setup
5.3 Results and Discussion
Table 1 compares the average runtimes and the number of iterations required for DPBVI and PBVI to converge on the MiniGrid DoorKey and Two-State environments. In both domains, DPBVI recovers the same optimal risk-neutral solution as PBVI, but the added cost of handling full return distributions significantly increases its runtime.
Table 1: Results of DPBVI and PBVI: average runtime and iteration counts. Both methods converge to the same value function.
@@ 5.3 Results and Discussion
In MiniGrid DoorKey, PBVI completes in $\approx 0.014$ seconds on average, whereas DPBVI requires $\approx 54.8$ seconds, though both converge in 11 iterations. For the Two-State problem, DPBVI is also slower ($0.3966$s vs. $0.0825$s), despite both methods converging in 789 iterations. These patterns illustrate how distributional operations can become computationally expensive, yet DPBVI preserves the optimal solution under risk-neutral objectives and paves the way for future risk-sensitive extensions.
6 Conclusion
In this paper, we presented the first comprehensive study of Distributional Reinforcement Learning in partially observable domains. Our main theoretical result showed that the partially observable distributional Bellman operator $\widetilde{\mathcal{T}}_{PO}^{\pi}$ and its optimality counterpart $\widetilde{\mathcal{T}}^{*}_{PO}$ are $\gamma$-contractions in the supremum p-Wasserstein metric ($1\leq p<\infty$). These results extend the well-known contraction properties of distributional Bellman operators in fully observable settings to POMDPs. As part of our analysis, we introduced $\psi$-vectors as the distributional analogs of $\alpha$-vectors, proving that under risk-neutral control, the optimal distributional value function in a POMDP can be represented by a finite collection of these $\psi$-vectors while preserving PWLC in the Wasserstein metric space. This representation underlines how a distributional perspective naturally generalizes classical POMDP theory.
Building on these foundations, we described DPBVI, a risk-neutral algorithm that applies the distributional Bellman optimality operator to a finite set of belief points. This extension parallels the classical PBVI procedure yet replaces scalar $\alpha$-vectors with distributional $\psi$-vectors, thereby capturing richer information about uncertainty in returns. We showed that DPBVI recovers the same solution as PBVI when taking expectations, confirming that the distributional approach does not compromise solution quality in the risk-neutral regime. While our primary focus was on risk-neutrality, this work provides a foundation for risk-sensitive approaches to partial observability, which may require new techniques and theoretical insights to handle more general utility functions and objectives. We hope these developments foster further exploration of DistRL in practical domains where risk and uncertainty are fundamental to decision-making.

### v2

@@ B.1 Environments
This environment, adapted from the MiniGrid suite (Chevalier-Boisvert et al., 2023, MiniGrid-DoorKey-5x5-v0), places an agent in a 5x5 grid containing a locked door and a key. The agent’s objective is to reach a goal state behind a locked door requiring a key. We assume both perfect sensor and transition models, a discount factor $\gamma=0.9$, a reward of $1.0$ upon reaching the goal state, and a known start state. Because the start state is known and the dynamics models are perfect, we set $\mathcal{B}$ to include complete confidence in being in one of the ten states along the optimal path (i.e., $|\mathcal{B}|=10$).
@@
B.2 Setup
We disable belief-point expansion in both environments and fix a finite set $\mathcal{B}$ of belief states as described in Section B.1. PBVI is implemented as in (Pineau et al., 2003), while DPBVI follows the backup procedure from Section 4. Return distributions in DPBVI use 51 categorical atoms, with each $\psi$-vector initialized to place all probability mass at zero.
@@ B.2 Setup
B.3 Results and Discussion
Table 1 compares the average runtimes and the number of iterations required for DPBVI and PBVI to converge on the MiniGrid DoorKey and Two-State environments. In both domains, DPBVI recovers the same optimal risk-neutral solution as PBVI, though with significantly higher runtime due to the additional cost of distributional operations.
Table 1: Results of DPBVI and PBVI: average runtime and iteration counts for convergence criterion $\epsilon=1\mathrm{e}{-3}$.
@@ B.3 Results and Discussion
To validate that DPBVI converges to the same value function as PBVI, we examine the maximum relative error between their value functions across belief points over successive backups. These results are shown in Figure 1 for both environments with a convergence criterion of $\epsilon=1\mathrm{e}{-3}$.
(a) Two-State Noisy-Sensor Environment
(b) MiniGrid DoorKey Environment
Figure 1: Maximum relative error between DPBVI and PBVI value functions per iteration with convergence criterion $\epsilon=1\mathrm{e}{-3}$.
In both domains, we observe very small value differences throughout the run, confirming that DPBVI effectively recovers the same solution as PBVI. In the Two-State environment (Figure 1(a)), the relative error increases gradually and exhibits intermittent spikes. We hypothesize these spikes arise from categorical projection error, floating-point precision limits, and bootstrapping. In MiniGrid DoorKey (Figure 1(b)), the relative error increases slightly before plateauing around $1.2\mathrm{e}{-7}$, again validating close alignment between both methods.
To test whether the value function discrepancy is a function of the number of backups—suggesting compounding error or a design flaw in DPBVI—we tightened the convergence threshold to $\epsilon=1\mathrm{e}{-6}$ and set the iteration limit to $10,000$ in the Two-State environment. Results are shown in Figure 2.
Figure 2: Maximum relative error between DPBVI and PBVI value functions with convergence threshold $\epsilon=1\mathrm{e}{-6}$ in the Two-State Noisy-Sensor environment.
As shown in Figure 2, the relative error increases over approximately the first $1{,}500$ backups, peaks just below $3\mathrm{e}{-4}$, and then gradually decreases and plateaus near $5\mathrm{e}{-5}$. This behavior supports our hypothesis that the discrepancy arises from bootstrapping, floating-point error, and categorical projection—rather than compounding error over time. Notably, neither PBVI nor DPBVI fully converged under this stricter threshold, yet the error remained bounded and decreased with additional backups. The MiniGrid DoorKey results remained unchanged under this more restrictive convergence criterion.